A recap of the ringdown measurements made at the 40m in mid 2012, the hardware that was installed for the same and results from the measurements.
A trans PD was installed (elog 7122).
This PD does not exist in the trans path anymore.
The polarity in the MC servo was flipped with the MC WFS disabled and the PD trans signal was used to look at the cavity ringdown.
Ringdown time = 13 us
Cavity finesse from the measurement = 453 (inconsistent with actual finesse).
Attachment: Ringdown measurement and fits
The AOM was installed before the PMC. The AOM was driven by the driver installed on the PSL table. (RF power ~1.5W @ 80MHz @1.0V modulation input). An RF switch was installed to control the AOM driver input. ZASWA-2-50DR+ was installed.
Note: The AOM was used by the ISS crew after this. So the RF switch has been removed and the AOM is no more a part of the ringdown setup.
No measurements were made as we could not obtain relevant TTL signals to control the switch remotely.